OpenStudy (anonymous):

What is the probability that two people selected at random have the same birthday? Ignore leap years.

OpenStudy (anonymous):

I think it's 1/133,225 but I'm not sure if I did it right.

OpenStudy (anonymous):

I need more information

OpenStudy (anonymous):

How did you get your answer?

OpenStudy (kropot72):

Given that the birthday of either of the two people is on a particular day, the probability that the birthday of the other person is on the same day is 1/365. Note that this is different from the situation where a day is selected at random and the probability that the two randomly selected people both have their birthdays on the same randomly selected day. In this case the probability is given by \[P(both\ have\ birthday\ on\ same\ randomly\ selected\ day)=\frac{1}{365} \times \frac{1}{365}=\frac{1}{133225}\]

OpenStudy (anonymous):

Well explained kropot72.

OpenStudy (kropot72):

@mrbarry Thank you :)